OpenStudy (anonymous):

Find the product. (-4·3·2)^2 is it -288?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Because : \((-)^2 = +\)

OpenStudy (anonymous):

576

OpenStudy (anonymous):

anyways its 576

OpenStudy (anonymous):

Yes it is 576..

OpenStudy (usukidoll):

order of operations PEMADAS (-4·3·2)^2 (-4)(3)(2) = -4(6) =-24 (-24)^2 --> write -24 two times like this (-24)(-24) a negative times a negative will be a positive number. 576